OpenStudy (anonymous):

Is this even odd or neither, and justify answer f(x)= [x+sin(x)]/cos(x) Is this even odd or neither, and justify answer f(x)= [x+sin(x)]/cos(x) @Mathematics

OpenStudy (anonymous):

it is odd because \[x+\sin(x)\] is odd and \[\cos(x)\] is even

OpenStudy (anonymous):

your teacher wants you to write \[f(x)=\frac{x+\sin(x)}{\cos(x)}\] \[f(-x)=\frac{-x+\sin(-x)}{\cos(-x)}=\frac{-x-\sin(x)}{\cos(x)}=-\frac{x+\sin(x)}{\cos(x)}=-f(x)\]

OpenStudy (anonymous):

lol my teacher uses this stuff in calc 3 all the time and we think he just does it to show off

OpenStudy (anonymous):

the second equal sign because \[\sin(-x)=-\sin(x)\text{ and }\cos(-x)=\cos(x)\]

OpenStudy (anonymous):

i guess it is useful for doing some integrals to know that the function is odd. also will tell you what power series expansion contains only odd exponents